The following table is a frequency table of the scores obtained in a competition. Use the table answer the questions below.

Classes

Frequency(f)

10 - 13 4

13 - 16 6

16 - 19 12

19 - 22 14

22 - 25 4

Total 40

a. Find the mean, median and mode of the score. [2,2,2]

b. Find the range, variance, and standard deviation. [1,3,1]

c. Find the coefficient of variation. [2]

d. Compute the interquartile range.


1
Expert's answer
2021-05-18T05:57:10-0400

a.

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 Class & f & m & d & f\cdot d & f\cdot d^2 & cf\\ \hline
10-13 & 4 & 11.5 &-2 &- 8 & 16 & 4 \\
13-16 & 6 & 14.5 & -1 & -6 & 6 & 10 \\
16-19 & 12 & 17.5 & 0 & 0 & 0 & 22 \\
19-22 & 14 & 20.5 & 1& 14 & 14 & 36 \\
22-25 & 4 & 23.5 & 2 & 8 & 16 & 40 \\
 
\end{array}"



"A=17.5"

"d=\dfrac{x-A}{h}, h=3"


"Mean=\bar{x}=A+\dfrac{\sum f d}{n}\cdot h"

"=17.5+\dfrac{8}{40}\cdot 3=18.1"

The median class is 16-19.

"L=16, n=40"

Cumulative frequency of the class preceding the median class "cf=10"

Frequency of the median class "f=12"

Class length of median class "c=3"

"Median\ M=L+\dfrac{\dfrac{n}{2}- cf}{f}\cdot c"

"=16+\dfrac{\dfrac{40}{2}- 10}{12}\cdot 3=18.5"


Maximum frequency is14.

The mode class is 19-22.

"L=19"

Cumulative frequency of the class preceding the median class "cf=10"

Frequency of the mode class "f_1=14"

Frequency of the preceding class "f_0=12"

Frequency of the succeeding class "f_2=4"

Class length of mode class "c=3"


"Z=L+\dfrac{f_1-f_0}{2f_1-f_0-f_2}\cdot c"

"=19+\dfrac{14-12}{2(14)-12-4}\cdot 3=19.5"

b.


"Range=\dfrac{22+25}{2}-\dfrac{10+13}{2}=12"


"S^2=\dfrac{\sum f\cdot d^2-\dfrac{(\sum f\cdot d)^2}{n}}{n-1}\cdot h^2"

"=\dfrac{52-\dfrac{(8)^2}{40}}{40-1}\cdot 3^2=\dfrac{151.2}{13}\approx11.63077"

"S=\sqrt{S^2}=\sqrt{\dfrac{151.2}{13}}\approx3.4104"

c.

"Coefficient \ of \ Variation=\dfrac{S}{\bar{x}}\cdot 100\%"

"=\dfrac{\sqrt{\dfrac{151.2}{13}}}{18.1}\cdot 100\%\approx18.842\%"

d.

Class with "(\dfrac{n}{4})^{th}"  value of the observation in "cf" column

"=(\dfrac{40}{4})^{th}" value of the observation in "cf" column

"=(10)^{th}" value of the observation in "cf" column

and it lies in the class 13-16.

"Q_1" class: "13-16"

"L=13"


"Q_1=L+\dfrac{\dfrac{n}{4}- cf}{f}\cdot c"

"=13+\dfrac{\dfrac{40}{4}- 4}{6}\cdot 3=16"

Class with "(\dfrac{3n}{4})^{th}"  value of the observation in "cf" column

"=(\dfrac{3\cdot 40}{4})^{th}" value of the observation in "cf" column

"=(30)^{th}" value of the observation in "cf" column

and it lies in the class 19-22.

"Q_3" class: "19-22"

"L=19"


"Q_3=L+\dfrac{\dfrac{3n}{4}- cf}{f}\cdot c"

"=19+\dfrac{\dfrac{3\cdot 40}{4}- 22}{14}\cdot 3\approx20.7143"

"IQR=Q_3-Q_1\approx20.7143-16\approx4.7143"